Suppose researchers perform a large-sample test of a population proportion where the null hypothesis is that the population proportion is 0.4, and the alternative hypothesis is that the population proportion is greater than 0.4 (one-sided). They obtain a z-statistic of 2 under the null hypothesis. What is the (one-sided) p-value?
Answer #1

P value = P ( X > 2 ) = 1 - P ( Z < 2 )  = 0.0228
looking for the value Z = 2 in standard normal table